# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ Abstract
Results. From our analysis of the stellar contributions to the continuum flux, we find that the semi-regular variables all have smaller physical radii and fainter monochromatic luminosities than the Mira variables. Comparing these properties with pulsation periods, we find a positive trend between stellar radius and period only for the Mira variables with periods above 300 days and a positive trend between the period and the monochromatic luminosity only for the red supergiants and the most extreme AGB stars with periods above 500 days. We find that the continuum emission at 1.24 mm can be classified into four groups. “Featureless” continuum emission is confined to the (unresolved) regions close to the star for five stars in our sample, relatively uniform extended flux is seen for four stars, tentative bipolar features are seen for three stars, and the remaining five stars have unique or unusual morphological features in their continuum maps. These features can be explained by binary companions to 10 out of the 14 AGB stars in our sample.
@@
At smaller angular scales, recent developments in adaptive optics have enabled dust imaging in the optical and near infrared at very high resolution. For example, Montargès et al. (2023) imaged a sample of 14 evolved stars (a subset of the sample we present in this work) at resolutions of mas. They presented maps of scattered polarised light which showed complex interactions in the inner circumstellar environments (within 10 to a few 100s of au). Even higher resolution imaging has been carried out for a few individual stars, spatially resolving their surfaces and showing non-uniform photospheres (Ohnaka et al., 2016; Paladini et al., 2018; Khouri et al., 2020), and convective time scales on the order of a month (Vlemmings et al., 2024). These findings are in agreement with high resolution 3D models of stellar convective envelopes, which show patchy dust formation (Höfner & Freytag, 2019) and dust forming in the wake of atmospheric shocks (Freytag & Höfner, 2023). High resolution ALMA observations of individual stars have also revealed inhomogeneities in the inner circumstellar envelope in the continuum and in molecular emission (Velilla-Prieto et al., 2023), further emphasising the non-uniform formation of dust at the surface of an AGB star.
@@ 2 Observations
SV Aqr | 23:22:45.4002 | 10:49:00.188 | 445 (65, 90) a | SRb * | 93, 231.8 | Extended, mid, compact | LABEL:fig:svaqr |
@@
Gru | 22:22:44.2696 | 45:56:53.006 | 164 (12, 10) b | SRb | 195.5, 128* | Extended, mid, compact, ACA | LABEL:fig:pi1gru |
@@
V PsA | 22:55:19.7228 | 29:36:45.038 | 299 (11, 12) b | SRb | 148, 105 | Extended, mid, compact | LABEL:fig:vpsa |
@@ 2.1 ATOMIUM sample
The ATOMIUM sample consists of 17 evolved stars, of which 14 are AGB stars and three are RSGs (VX Sgr, AH Sco and KW Sgr). The sample was constructed so as to cover a range of mass-loss rates and variability types (e.g. Mira variables and semiregular variables). Of the AGB stars, 12 are oxygen-rich (M-type, C/O ) and two are S-type stars with C/O (W Aql and Gru). Coincidentally, the two S-type stars are also known to have binary companions (Ramstedt et al., 2011; Mayer et al., 2014), though Gru has been proven to be a triple system (Homan et al. 2020; Montargès et al. 2025; Esseldeurs et al. subm.). The AGB star R Hya also has a companion star, but at a projected separation of au (R Hya B, Smak, 1964; Kervella et al., 2022, aka Gaia DR3 6195030801634430336), it is too distant to have a significant impact on the circumstellar envelope.
@@ 2.1.2 Variability
We list periods and variability types in Table 1 for all the stars in our sample. The variability types and periods were taken from the International Variable Star Index (VSX222VSX: www.aavso.org/vsx/index.php). These periods were also further verified by cross-checking the VSX period with light curves from the American Association of Variable Star Observers (AAVSO333AAVSO: www.aavso.org), All Sky Automated Survey (ASAS, Pojmanski, 2002) and the ASAS-SN Variable Stars Database (Shappee et al., 2014; Jayasinghe et al., 2019) using the Lomb-Scargle periodogram (VanderPlas, 2018), as the time series are unevenly sampled. Among the AGB stars we have a mix of Mira variables and semi-regular (SR) variables of types SRa and SRb. The RSGs are all SRc, as expected. The SRb stars have a secondary period as well as a primary period listed in Table 1.
In Appendix A we discuss individual stellar periods in more detail. To briefly summarise the key points, SV Aqr was reported in earlier ATOMIUM papers as a long period variable (e.g. Gottlieb et al., 2022). For the present work, we looked further into the pulsations of SV Aqr and were able to define a primary period of 93 days, as described in Appendix A. For Gru, we identified an additional long secondary period (of years), also described in Appendix A. R Aql and R Hya have been reported to have shortening periods (Wood & Zarro, 1981; Greaves & Howarth, 2000; Zijlstra et al., 2002; Joyce et al., 2024), possibly owing to undergoing thermal pulses in the recent past (a few thousand or so years ago). This is also discussed further in Sect. 5.3.
@@ 2.2 ALMA 12m Array
Observations were taken with three configurations of the ALMA 12m Array in Band 6 over the frequency range 214–270 GHz as part of the ATOMIUM Large Programme (2018.1.00659.L, PI: L. Decin). Typical resolutions for the extended, mid and compact array configurations are 25 mas, 200 mas, and 1″. A detailed discussion of the data reduction is given in Gottlieb et al. (2022) and a complete list of the spectral line IDs and their properties is given in Wallström et al. (2024). The maximum recoverable scales (MRS) are – for the extended array data, for the mid array data, and 8– for the compact array data. This is discussed further in Sect. 4.1 and more precise MRS for each star and configuration are given in Table E.2 of Gottlieb et al. (2022). In this paper we focus on the continuum data, which we now analyse in detail. However, the basic parameters of the individual array continuum observations, including observation dates, are given in Table E.2 of Gottlieb et al. (2022).
@@
The ALMA 12m array data taken in the compact configuration (lowest resolution) were observed for the shortest time spans and so have lower signal to noise (S/N) and positional precision. For each star and array configuration, we identified and excluded spectral lines and self-calibrated the phase-referenced continuum data, as described in Gottlieb et al. (2022). The flux scale of individual observations should be accurate to 5–7% at Band 6 (Cortes et al., 2023) and our observations met this. Where there were discrepancies, we took the mid configuration flux densities as the most reliable standard to rescale the other configurations (based on baseline lengths in common). The resulting flux scale is nominally accurate to although detailed examination (e.g. of R Hya by Homan et al., 2021) suggests that it may be better.
@@
We fitted 2-D Gaussian components to the continuum images after applying phase-reference solutions only. The accuracy of Gaussian fitting is limited by (beam size)/(S/N), where S/N is 50 in all cases; the main astrometric uncertainty comes from transfer of calibration (Gottlieb et al., 2022). The most accurate astrometric positions (2 mas) were obtained for extended configuration images. The mid and compact array data were adjusted to the position of the extended array data before combination. Thus the stellar reference positions are for the epoch 2019 June 23 – 2019 July 12 (observing dates for each star are given individually in Table E1 of Gottlieb et al., 2022).
@@
When we discuss combined data, we are referring to the ALMA 12m array data from the extended, mid and (for most stars) compact configurations that have been combined to improve overall sensitivity while maintaining a high angular resolution. The observed data for each separate array observation were weighted during pipeline calibration according to the integration times, channel widths, and amplitude calibration variance, but no further weighting was done in self-calibration (and flagging was done after channel averaging). The configurations were combined with equal weight.
@@
For any stars with signs of artefacts such as negatives exceeding , we re-examined the data carefully. Calibration and flux scale errors are symmetric (or anti-symmetric for phase) in the stellar image, whilst angular misalignments tend to show up as anti-symmetric central adjacent errors after UD subtraction. In a few cases we repeated the self-calibration of individual configurations with different solution intervals and other constraints to optimise the continuum. However, the main cause of artefacts was usually misalignment of the flux or position between configurations, since even within the ALMA tolerances an e.g. 5% flux scale error exceeds the dynamic range for the brighter stars. We therefore aligned the configurations to minimise the relative errors.
@@ 2.3 ALMA Compact Array
A subset of the ATOMIUM sample was also observed in Band 3 with the ALMA Compact Array (Morita Array; project 2019.1.00187.S, PI: T. Danilovich). The observations were taken over discrete spectral windows in the frequency range 97–113 GHz. The central frequency for the continuum images is 104.96 GHz. The MRS for these observations fall in the range 64–77″. The flux scale accuracy of this data is around 5% (Cortes et al., 2023). Note that we refer to observations with the ALMA Compact Array as “ACA data” and to observations with the compact configuration of the ALMA 12m Array as “compact data”. The details of the ACA data are given in Table 9. In Fig. 12 we show all the continuum images obtained using the ACA. We also present some line observations obtained concurrently with the ACA continuum in Appendix D.
All the ACA data used comes from the standard ALMA pipeline aside from RW Sco, for which we created a custom dataset that was not primary beam corrected. The reason for this is a confounding nearby source, approximately south of RW Sco, which is intrinsically brighter than the AGB star once the primary beam correction is applied. This is discussed further in Appendix C.
@@ 2.4 Fitting uniform discs
To estimate the stellar sizes and flux density, we fit a Uniform Disc (UD) model to the calibrated, line-free combined Band 6 visibilities for each source. We tested Gaussian fits but found higher residuals indicating worse fits. Trying to fit a UD + a point source at the stellar position also gave worse results, including relatively bright residuals within the stellar disc diameter. This suggests that any stellar-related flux not fit by a UD is not concentrated in a single point and is not symmetric, i.e. has a random or more complex and blotchy distribution, which cannot be fit in the uv plane. We also found that the irregularities in the sources are greater than any differences between a uniform disc or a limb-darkened disc (where a limb-darkened disc is similar to what Bojnordi Arbab et al. (2024) found from their models).
@@
For the UD fits we used the CASA task uvmodelfit with the following exceptions. In the case of AH Sco, RW Sco and Gru, we used the uvmultifit package (Martí-Vidal et al., 2014). For AH Sco and IRC+10011, we used the extended configuration data only as there are irregularities on short spacings; for RW Sco we forced the position to be at the phase centre, as it has very weak continuum. In each case, the starting model was the optical diameter and peak flux density from Tables 1 and E.2 in Gottlieb et al. (2022). All parameters (peak, position, diameter, ellipticity, position angle) were allowed to vary except for the special case of Gru, for which a close companion ( Gru C, separation ) was reported by Homan et al. (2021). Hence, for Gru we first subtracted the companion and then fixed the position and ellipticity (none) to avoid contamination by the companion residuals. This two-component fit was done using only the extended array data, but we find the UD model agrees with the full dataset, except at the shortest baselines (see Fig. 13). Gru C is too faint and compact to be resolved even by component fitting so we do not give its size (see Fig. LABEL:fig:pi1gru, in which it is comparable to the extended data beam size). We determine an integrated flux density for Gru C of 2.78 mJy.
@@
For SV Aqr, attempts to fit to the visibilities for all configurations failed to resolve the star and could only measure the flux density within the central 15 mas with any confidence. The result of this fit is shown in Fig. 13. However, to retrieve more reliable stellar parameters, we performed a fit using the extended configuration data only, which retrieved the parameters with S/N . It is these parameters that are included in Table 2.
@@
The fitting routines report the uncertainties in the fits. The disc fits were almost circular ( minor/major axis ratio ¿0.9) for most sources, except RW Sco and KW Sgr (which have low SNR) and IRC10529 (), and U Her (), for which the diameter listed in Table 2 is the longer axis. We assume that the stars are unlikely to be elliptical to this degree, but instead that deviations are either due to elliptical beams (Table 8) or stellar surface irregularities probably distributed randomly. We thus take the ellipticity as a contribution to the uncertainty in the fit. The other contributions are the observational errors. Phase errors contribute a size uncertainty proportional to the beam size/SNR, and the flux density errors are mostly dominated by the flux scale uncertainty except for the weakest sources where the noise is significant (see Gottlieb et al., 2022).
@@
We were unable to find good fits for KW Sgr or RW Sco because these two sources have low SNR and small angular sizes, consistent with being unresolved. (Note that the other stars, aside from SV Aqr, have S/N ¿ 50.) We still list the best fitting parameters for these stars in Table 2 and show the fits in Fig. 13 but note the very large uncertainties on the UD diameters. In Table 1, we give the right ascension and declination (ICRS) determined from the UD fits, which in all cases are within 1 mas of the astrometric positions (see Sect. 2.2 and Gottlieb et al. 2022). Preliminary UD sizes were reported for 5 of the ATOMIUM sources previously (R Aql, U Her, R Hya, T Mic, and S Pav, Baudry et al., 2023, see also Homan et al. 2021 for R Hya). Those sizes were within 1% of what we find in the present work, except for R Hya where the size difference is .
Based on our UD fits we also calculated brightness temperatures for our sample, using 241.75 GHz as the central frequency of our combined data. An explanation of this calculation is given in Appendix E. The resultant brightness temperatures are given in Table 2, excluding RW Sco, for which the UD fit values were too uncertain to implement. The uncertainties we find for KW Sgr and SV Aqr are very large, but we list the obtained values in the table for completion. The flux densities of these sources are the lowest in our sample. Note that KW Sgr is our most distant target. For the sample overall, the derived brightness temperatures are in the range 760–2300 K, all systematically cooler than the effective stellar temperatures collected from the literature (see Table 3), but in line with the earlier results of Vlemmings et al. (2019).
We subtracted the UD models from the combined visibility data and imaged the residuals, which we show alongside the continuum images in Figures 2–LABEL:fig:ahsco. We weighted the baselines during imaging (see Sect. 3) to provide a larger synthesised beam than was used for the un-subtracted combined images, to give better sensitivity to features with low surface brightness. The exceptions are GY Aql and Gru where we use a synthesised beam of comparable size to the combined image in order to resolve the faint, but compact and distinct, offset emission. The details of the UD subtracted images are given in Table 10. For SV Aqr we show the residual image after subtracting the original UD fit to the combined data, which represents contributions from the star and possibly some inner dust. We caution that the resulting features are not as reliable as for the other stars and should not be over-interpreted.

### v2

# ATOMIUM: Continuum emission and evidence of dust enhancement from binary motion

@@ Abstract
Results. From our analysis of the stellar contributions to the continuum flux, we find that the semi-regular variables all have smaller physical radii and fainter monochromatic luminosities than the Mira variables. Comparing these properties with pulsation periods, we find a positive trend between stellar radius and period only for the Mira variables with periods above 300 days and a positive trend between the period and the monochromatic luminosity only for the red supergiants and the most extreme AGB stars with periods above 500 days. We find that the continuum emission at 1.24 mm can be classified into four groups. “Featureless” continuum emission is confined to the (unresolved) regions close to the star for five stars in our sample, relatively uniform extended flux is seen for four stars, tentative elongated features are seen for three stars, and the remaining five stars have unique or unusual morphological features in their continuum maps. These features can be explained by binary companions to 10 out of the 14 AGB stars in our sample.
@@
At smaller angular scales, recent developments in adaptive optics have enabled dust imaging in the optical and near infrared at very high resolution. For example, Montargès et al. (2023) imaged a sample of 14 evolved stars (a subset of the sample we present in this work) at resolutions of mas. They presented maps of scattered polarised light which showed complex interactions in the inner circumstellar environments (within 10 to a few 100s of au). Even higher resolution imaging has been carried out for a few individual stars, spatially resolving their surfaces and showing non-uniform photospheres (Ohnaka et al., 2016; Paladini et al., 2018; Khouri et al., 2020), and convective time scales on the order of a month for at least one star (R Dor, Vlemmings et al., 2024). These findings are in agreement with high resolution 3D models of stellar convective envelopes, which show patchy dust formation (Höfner & Freytag, 2019) and dust forming in the wake of atmospheric shocks (Freytag & Höfner, 2023). High resolution ALMA observations of individual stars have also revealed inhomogeneities in the inner circumstellar envelope in the continuum and in molecular emission (Velilla-Prieto et al., 2023; Baudry et al., 2023; Gottlieb et al., 2022), further emphasising the non-uniform formation of dust at the surface of an AGB star.
@@ 2 Observations
| SV Aqr | 23:22:45.4002 | 10:49:00.188 | 445 (65, 90) a | SRb * | 93, 231.8* | Extended, mid, compact | LABEL:fig:svaqr |
@@
| Gru | 22:22:44.2696 | 45:56:53.006 | 164 (12, 10) b | SRb | 195.5, 5750* | Extended, mid, compact, ACA | LABEL:fig:pi1gru |
@@
| V PsA | 22:55:19.7228 | 29:36:45.038 | 299 (11, 12) b | SRb | 148* | Extended, mid, compact | LABEL:fig:vpsa |
@@ 2.1 ATOMIUM sample
The ATOMIUM sample consists of 17 evolved stars, of which 14 are AGB stars and three are RSGs (VX Sgr, AH Sco and KW Sgr). The sample was constructed so as to cover a range of mass-loss rates and variability types (e.g. Mira variables and semiregular variables). Of the AGB stars, 12 are oxygen-rich (M-type, C/O ) and two are S-type stars with C/O (W Aql and Gru). Coincidentally, the two S-type stars are also known to have binary companions (Ramstedt et al., 2011; Mayer et al., 2014), though Gru has been proven to be a triple system (Homan et al., 2020; Montargès et al., 2025; Esseldeurs et al., 2025). The AGB star R Hya also has a companion star, but at a projected separation of au (R Hya B, Smak, 1964; Kervella et al., 2022, aka Gaia DR3 6195030801634430336), it is too distant to have a significant impact on the circumstellar envelope.
@@ 2.1.2 Variability
We list periods and variability types in Table 1 for all the stars in our sample. The variability types and periods were taken from the International Variable Star Index (VSX222VSX: www.aavso.org/vsx/index.php). These periods were also further verified by cross-checking the VSX period with light curves from the American Association of Variable Star Observers (AAVSO333AAVSO: www.aavso.org), All Sky Automated Survey (ASAS, Pojmanski, 2002) and the ASAS-SN Variable Stars Database (Shappee et al., 2014; Jayasinghe et al., 2019) using the Lomb-Scargle periodogram (VanderPlas, 2018), as the time series are unevenly sampled. Among the AGB stars we have a mix of Mira variables and semi-regular (SR) variables of types SRa and SRb. The RSGs are all SRc, as expected. The SRb stars aside from V PsA have a secondary period as well as a primary period listed in Table 1.
In Appendix A.1 we discuss individual stellar periods in more detail. To briefly summarise the key points, SV Aqr was reported in earlier ATOMIUM papers as a long period variable (e.g. Gottlieb et al., 2022). For the present work, we looked further into the pulsations of SV Aqr and were able to define a primary period of 93 days, as described in Appendix A.1. For Gru, we identified a long secondary period (LSP, of years), also described in Appendix A.1. R Aql and R Hya have been reported to have shortening periods (Wood & Zarro, 1981; Greaves & Howarth, 2000; Zijlstra et al., 2002; Joyce et al., 2024), possibly owing to undergoing thermal pulses in the recent past (a few hundred or so years ago). This is also discussed further in Sect. A.2.
@@ 2.2 ALMA 12m Array
Observations were taken with three configurations of the ALMA 12m Array in Band 6 over the frequency range 214–270 GHz as part of the ATOMIUM Large Programme (2018.1.00659.L, PI: L. Decin). Typical resolutions for the extended, mid and compact array configurations are 25 mas, 200 mas, and 1″. In Table 1 we note which stars were observed with which configurations, and give a reference to which figure their continuum images are shown in (Fig.s 2–LABEL:fig:ahsco). All the compact configuration images are shown in Fig. 14. A detailed discussion of the data reduction, including self-calibration, is given in Gottlieb et al. (2022) and a complete list of the spectral line IDs and their properties is given in Wallström et al. (2024). The maximum recoverable scales (MRS, the scale at which smooth regions of flux can be recovered with confidence) are – for the extended array data, for the mid array data, and 8– for the compact array data. This is discussed further in Sect. 3.1 and more precise MRS for each star and configuration are given in Table E.2 of Gottlieb et al. (2022). In this paper we focus on the continuum data, which we now analyse in detail. However, the basic parameters of the individual array continuum observations, including observation dates, are given in Table E.2 of Gottlieb et al. (2022). In Tables 7, 8, and 9 we give an overview of the continua observed with the extended, mid and compact array configurations.
@@
The ALMA 12m array data taken in the compact configuration (lowest resolution) were observed for the shortest time spans and so have lower signal to noise (S/N) and positional precision. For each star and array configuration, we identified and excluded spectral lines and self-calibrated the phase-referenced continuum data, as described in Gottlieb et al. (2022). The flux scale of individual observations should be accurate to 5–7% at Band 6 (Cortes et al., 2023) and our observations met this. Where there were discrepancies due to calibration uncertainty, we rescaled an individual configuration if it was inconsistent or, if all configurations disagreed, we took the mid configuration flux densities as the most reliable standard to rescale the other configurations. This flux comparison was done for baseline lengths common between configurations, hence was not affected by missing spacings. The resulting flux scale is nominally accurate to although detailed examination (e.g. of R Hya by Homan et al., 2021) suggests that it may be better.
@@
We fitted 2-D Gaussian components to the continuum images after applying phase-reference solutions only. The accuracy of Gaussian fitting is limited by (beam size)/(S/N), where S/N is 50 in all cases; the main astrometric uncertainty comes from transfer of calibration (Gottlieb et al., 2022). The most accurate astrometric positions (2 mas) were obtained for extended configuration images. The mid and compact array data were adjusted to the position of the extended array data before combination. Thus the stellar reference positions are for the epoch 2019 June 23 – 2019 July 12 (observing dates for each star are given individually in Table E1 of Gottlieb et al., 2022). Overall, data for each star was taken over a period of 6–12 months and the proper motions for each star range from 1.3–69 mas yr-1 (Gaia Collaboration et al., 2023), likely accounting for the most significant positional shifts. The only exception is KW Sgr, for which one set of mid observations were taken in May 2021, giving a total observation range of 30 months. The proper motion shift was minimal, however, since this distant star has the smallest proper motion of our sample.
@@
When we discuss combined data, we are referring to the ALMA 12m array data from the extended, mid and (for most stars) compact configurations that have been combined to improve overall sensitivity while maintaining a high angular resolution. An overview of the combined data is given in Table 10. The observed data for each separate array observation were weighted during pipeline calibration according to the integration times, channel widths, and amplitude calibration variance, but no further weighting was done in self-calibration (and flagging was done after channel averaging). The configurations were combined with equal weight. The combined MRS is generally as good as for the compact configuration data, except for AH Sco and KW Sgr, which were not observed with the compact configuration, and for which the MRS is as good as for the mid configuration data.
@@
For any combined images with signs of artefacts such as negatives exceeding , we re-examined the data carefully. Calibration and flux scale errors are symmetric (or anti-symmetric for phase) in the stellar image, whilst angular misalignments tend to show up as anti-symmetric central adjacent errors after subtracting a uniform disc (see Sect. 2.4). In a few cases we repeated the self-calibration of individual configurations with different solution intervals and other constraints to optimise the continuum. However, the main cause of artefacts was usually misalignment of the flux or position between configurations, since even within the ALMA tolerances an e.g. 5% flux scale error exceeds the dynamic range for the brighter stars. We therefore aligned the configurations to minimise the relative errors.
@@ 2.3 ALMA Compact Array
A subset of the ATOMIUM sample was also observed in Band 3 with the Morita Array (ALMA Compact Array; project 2019.1.00187.S, PI: T. Danilovich). The observations were taken over discrete spectral windows in the frequency range 97–113 GHz. The central frequency for the continuum images is 104.96 GHz. The MRS for these observations fall in the range 64–77″. The flux scale accuracy of this data is around 5% (Cortes et al., 2023). Note that we refer to observations with the Morita Array as “ACA data” and to observations with the compact configuration of the ALMA 12m Array as “compact data”. The details of the ACA data are given in Table 11. In Fig. 15 we show all the continuum images obtained using the ACA.
All the ACA data used comes from the standard ALMA pipeline (and has not been self-calibrated) aside from RW Sco, for which we created a custom dataset that was not primary beam corrected. The reason for this is a confounding nearby source, approximately south of RW Sco, which is intrinsically brighter than the AGB star once the primary beam correction is applied. This is discussed further in Appendix C. We present some line observations obtained concurrently with the ACA continuum in Appendix D.
@@ 2.4 Fitting uniform discs
For the Band 6 observations of each star, we show the amplitudes of the visibilities as a function of distance in Fig. 16. These plots generally indicate a compact component and, in many cases, some extended features, the latter discussed in more detail in Sect. 3. At radio frequencies, e.g. wavelengths of cm to sub-mm, the main compact contribution to the continuum emission is expected to come from electron-neutral free-free emission in an extended radio photosphere, above the stellar photosphere (Reid & Menten, 1997; Matthews et al., 2015; Vlemmings et al., 2019). Emission offset from the stellar position, assumed to be the continuum peak, most likely originates from dust surrounding the stellar photosphere and radiosphere (see Sect. 3).
@@
To estimate the stellar sizes and flux densities, we fit a Uniform Disc (UD) model to the calibrated, line-free combined Band 6 visibilities for each source. We tested Gaussian fits but found higher residuals indicating worse fits. Trying to fit a UD + a point source at the stellar position also gave worse results, including relatively bright residuals within the stellar disc diameter. This suggests that any stellar-related flux not fit by a UD is not concentrated in a single point and is not symmetric, i.e. has a random or more complex and blotchy distribution, which cannot be fit in the uv plane. We also found that the irregularities in the sources are greater than any differences between a uniform disc or a limb-darkened disc (where a limb-darkened disc is similar to what Bojnordi Arbab et al. (2024) found from their models).
@@
For the UD fits we used the CASA task uvmodelfit to fit a single UD to the data. In the case of AH Sco, RW Sco and Gru, we used the uvmultifit package (Martí-Vidal et al., 2014), which has many capabilities, including the ability to fit multiple components and more detailed constraints at low S/N, which allowed us to ensure that the diameter and flux did not simply match pre-set limits. For our sample of evolved stars, we report only single component fits as adding more components, whether points or extended emission, did not converge to a stable solution, suggesting that any star-spots and/or extended dust cannot be described by simple, symmetric components. For AH Sco and IRC+10011, we used the extended configuration data only as there are irregularities on short spacings. These are probably due to large-scale flux which is irregularly or asymmetrically distributed, possibly simply contributing a small excess on short baselines but smooth enough on large scales to be resolved out for long baselines. For most stars such irregularities did not affect the fit, but for AH Sco (because of low S/N) and IRC+10011 (significant resolved-out flux, see Sect. 3.1), the errors in the fit were much greater if short baselines were included. For RW Sco we forced the position to be at the phase centre, as it has very weak continuum. In each case, the starting model was the literature optical diameter and peak flux density from Tables 1 and E.2 in Gottlieb et al. (2022). All parameters (peak, position, diameter, ellipticity, position angle) were allowed to vary except for the special case of Gru, for which a close companion ( Gru C, separation ) was reported by Homan et al. (2021). Hence, for Gru we first subtracted the companion and then fixed the position and ellipticity (none) to avoid contamination by the companion residuals. This two-component fit was done using only the extended array data, but we find the UD model agrees with the full dataset, except at the shortest baselines (see Fig. 16). Gru C is too faint and compact to be resolved even by component fitting so we do not give its size (see Fig. LABEL:fig:pi1gru, in which it is comparable to the extended data beam size). We determine an integrated flux density for Gru C of 2.78 mJy.
@@
For SV Aqr, attempts to fit to the visibilities for all configurations failed to resolve the star and could only measure the flux density within the central 15 mas with any confidence. The result of this fit is shown in Fig. 16. However, to retrieve more reliable stellar parameters, we performed a fit in the image plane using the extended configuration data only, which retrieved the parameters with S/N . It is these parameters that are included in Table 2.
@@
The fitting routines report the uncertainties in the fits. The disc fits were almost circular ( minor/major axis ratio ¿0.9) for most sources, except RW Sco and KW Sgr (which have low S/N) and IRC10529 (), and U Her (), for which the diameter listed in Table 2 is the longer axis. We assume that the stars are unlikely to be elliptical to this degree, but instead that deviations are either due to elliptical beams (Table 10) or stellar surface irregularities probably distributed randomly. We thus take the ellipticity as a contribution to the uncertainty in the fit. The other contributions are the observational errors. Phase errors contribute a size uncertainty proportional to the (beam size)/(S/N), and the flux density errors are mostly dominated by the flux scale uncertainty except for the weakest sources where the noise is significant (see Gottlieb et al., 2022).
@@
We were unable to find good fits for KW Sgr or RW Sco because these two sources have low S/N and small angular sizes, consistent with being unresolved. (Note that the other stars, aside from SV Aqr, have S/N ¿ 50.) We still list the best fitting parameters for these stars in Table 2 and show the fits in Fig. 16 but note the very large uncertainties on the UD diameters. In Table 1, we give the right ascension and declination (ICRS) determined from the UD fits, which in all cases are within 1 mas of the astrometric positions (see Sect. 2.2 and Gottlieb et al. 2022). Preliminary UD sizes were reported for 5 of the ATOMIUM sources previously (R Aql, U Her, R Hya, T Mic, and S Pav, Baudry et al., 2023, see also Homan et al. 2021 for R Hya). Those sizes were within 1% of what we find in the present work, except for R Hya where the size difference is . The improvement in the R Hya fit comes from the adjustments to the relative positions and fluxes described in Sect. 2.2, which enabled us to improve the dynamic range from to 6500.
Based on our UD fits we also calculated brightness temperatures for our sample, using 241.75 GHz as the central frequency of our combined data. An explanation of this calculation is given in Appendix E. The resultant brightness temperatures are given in Table 2, excluding RW Sco, for which the UD fit values were too uncertain to implement. The uncertainties we find for KW Sgr and SV Aqr are very large, but we list the obtained values in the table for completion. The flux densities of these sources are the lowest in our sample. Note that KW Sgr is our most distant target. For the sample overall, the derived brightness temperatures are in the range 760–2300 K, all systematically cooler than the effective stellar temperatures collected from the literature (see Table 4), but many are largely in line with the earlier results of Vlemmings et al. (2019), for four nearby AGB stars. The theoretical predictions of Bojnordi Arbab et al. (2024) give phase-dependent brightness temperatures K, depending on the model. Where our brightness temperatures are notably lower than this, we assume it is because newly-formed dust, located close to the surface of the star, has been captured by our best-fitting UD.
We subtracted the UD models from the combined visibility data and imaged the residuals, which we show alongside the continuum images in Figures 2–LABEL:fig:ahsco. We weighted the baselines during imaging (see Sect. 3) to provide a larger synthesised beam than was used for the un-subtracted combined images, to give better sensitivity to features with low surface brightness. The exceptions are GY Aql and Gru where we use a synthesised beam of comparable size to the combined image in order to resolve the faint, but compact and distinct, offset emission. The details of the UD subtracted images are given in Table 12. For SV Aqr we show the residual image after subtracting the original UD fit to the combined data (Fig. LABEL:fig:svaqr), which represents contributions from the star and possibly some inner dust. We caution that the resulting features are not as reliable as for the other stars and should not be over-interpreted.